# ISAC Network Planning: Sensing Coverage Analysis and 3-D BS Deployment Optimization

### v1

@@ IV-B2 Problem Transformation
$=$ $a_{1}\mathbf{v}^{T}_{n,k}{\mathbf{M}}^{-1}\mathbf{v}_{n,k}\!+\!\text{Tr}\left({\bf{q}}_{1}^{T}\mathbf{v}_{n,k}\right)\!+\!\text{Tr}\left(\mathbf{G}_{n,k}(\mathbf{U}_{n,k}^{c})^{T}{\mathbf{M}}^{-1}\mathbf{U}_{n,k}^{c}\right)\!,\vspace{-1.5mm}$
where $a_{1}=\mathbf{1}^{T}\mathbf{G}_{n,k}\mathbf{1}+\mathbf{e_{1}}^{T}\mathbf{G}_{n,k}\mathbf{1}+\mathbf{1}^{T}\mathbf{G}_{n,k}\mathbf{e_{1}}+\mathbf{e}^{T}\mathbf{G}_{n,k}\mathbf{e_{1}}$, and ${\bf{q}}_{1}^{T}=\left({\mathbf{M}}^{-1}\mathbf{U}_{n,k}^{c}\mathbf{G}_{n,k}\mathbf{1}\right)^{T}+\mathbf{1}^{T}\mathbf{G}_{n,k}(\mathbf{U}^{c}_{n,k})^{T}{\mathbf{M}}^{-1}+\mathbf{e_{1}}^{T}\mathbf{G}_{n,k}(\mathbf{U}_{n,k}^{c})^{T}{\mathbf{M}}^{-1}+\left({\mathbf{M}}^{-1}\mathbf{U}_{n,k}^{c}\mathbf{G}_{n,k}{\mathbf{e}}_{1}\right)^{T}$.
@@
$+a_{1}\lambda_{\max}\left({\mathbf{M}}^{-1}\right)+a_{1}\left(\mathbf{v}_{n,k}^{(r)}\right)^{T}{\mathbf{\tilde{M}}}\mathbf{v}_{n,k}^{(r)},\vspace{-1.5mm}$
@@
where $\tilde{\mathbf{M}}={\mathbf{M}}^{-1}-\lambda_{\max}\left({\mathbf{M}}^{-1}\right){\mathbf{I}}$, and $\lambda_{\max}({\mathbf{M}}^{-1})$ denotes the maximum eigenvalue of matrix ${\mathbf{M}}^{-1}$. In addition, in (34), the term $\mathbf{P}_{n,k}^{T}\mathbf{U}_{n,k}$ can be reduced by discarding any components independent of the optimization variable ${\bf{b}}_{n}$, leaving only the contributions $({{{\mathbf{1}}^{T}}{{\mathbf{P}}_{n,k}^{T}}+{{\mathbf{e}}^{T}_{1}}{{\mathbf{P}}_{n,k}^{T}}})\mathbf{v}_{n,k}$. By omitting the constant terms, the objective function can be approximated as
@@
$\mathop{\mathrm{min}}\limits_{{\mathbf{b}}_{n}}\frac{1}{K}\sum\limits_{k=1}^{K}\left(a_{2}\delta^{2}_{n,k}+a_{3}\delta_{n,k}+{\mathbf{q}}_{2}^{T}\frac{\mathbf{t}_{k}-{\mathbf{b}}_{n}}{\left\|\mathbf{t}_{k}-{\mathbf{b}}_{n}\right\|}\right),\vspace{-1.5mm}$ (37)
@@
where ${\mathbf{q}}_{2}^{T}=2a_{1}\left(\tilde{\mathbf{M}}\mathbf{v}_{n,k}^{(r)}\right)^{T}+{\mathbf{q}}_{1}^{T}-{{{\mathbf{1}}^{T}}{{\mathbf{P}}_{n,k}^{T}}-{{\mathbf{e}}^{T}_{1}}{{\mathbf{P}}_{n,k}^{T}}}$, and $a_{2}={\mathbf{G}}_{n,k}{[n,n]}$, $a_{3}=\sum\limits_{m=1,m\neq n}^{N}{\mathbf{G}}_{n,k}{[m,m]}$ $\left\|\mathbf{t}_{k}-{\mathbf{b}}_{m}\right\|^{\beta}$. Here, (37) is still non-convex due to $\frac{\mathbf{t}_{k}-{\mathbf{b}}_{n}}{\left\|\mathbf{t}_{k}-{\mathbf{b}}_{n}\right\|}$. By utilizing the trust region algorithm of [36], we introduce a constraint $\|{\mathbf{b}}_{n}-{\mathbf{b}}_{n}^{(r)}\|\leq\epsilon$ to limit the update region. In the trust region method, a first-order approximation is used at each iteration for constructing a local linear model of the objective function based on its gradient. In practice, this leads to computing the so-called Cauchy point [36], which is obtained by taking a step in the negative gradient’s direction scaled to the boundary of the trust region. Specifically, we adopt Taylor expansion to approximate the objective function as follow.
$\frac{\mathbf{t}_{k}-{\mathbf{b}}_{n}}{\left\|\mathbf{t}_{k}-{\mathbf{b}}_{n}\right\|}\approx\frac{\mathbf{t}_{k}-{\mathbf{b}}_{n}^{(r)}}{\|\mathbf{t}_{k}-{\mathbf{b}}_{n}^{(r)}\|}+g^{\prime}({\mathbf{b}}_{n})({\mathbf{b}}_{n}-{\mathbf{b}}_{n}^{(r)}),\vspace{-1.5mm}$ (38)
@@
where $g^{\prime}(\mathbf{b}_{n})=-\frac{1}{\|\mathbf{t}_{k}-\mathbf{b}_{n}^{(r)}\|}\,\Biggl{(}\,\mathbf{I}-\frac{(\mathbf{t}_{k}-\mathbf{b}_{n}^{(r)})(\mathbf{t}_{k}-\mathbf{b}_{n}^{(r)})^{T}}{\|\mathbf{t}_{k}-\mathbf{b}_{n}^{(r)}\|^{2}}\Biggr{)}$. Following the transformation above, every term in the objective function involving the decision variables is either a convex quadratic or a linear function. Then, for the communication constraint, we introduce auxiliary variables:
$z_{n,j}\geq\|{\mathbf{b}}_{n}-{\mathbf{u}}_{j}\|^{\alpha},\quad\forall j.\vspace{-1.5mm}$ (39)
@@
We define $C_{j}=1+\frac{1}{\sigma_{c}^{2}}\sum_{m\neq n}g_{m}\|\mathbf{b}_{m}-\mathbf{u}_{j}\|^{-\alpha}$, and apply the first-order Taylor expansion to approximate the communication rate of the user located at ${\bf{u}}_{j}$ as ${R_{c}\left({\bf{u}}_{j}\right)}\geq\log_{2}\left(C_{j}+\frac{g_{n}}{\sigma_{c}^{2}z_{n,j}^{(r)}}\right)-\frac{{g_{n}}{}}{\sigma_{c}^{2}(z_{n,j}^{(r)})^{2}\left(C_{j}+\frac{g_{n}}{\sigma_{c}^{2}z_{n,j}^{(r)}}\right)\ln 2}\left(z_{n,j}-z_{n,j}^{(r)}\right)=\tilde{R}_{j}$. Thus, the original non-convex communication constraint can be replaced by the following convex constraint:
$\frac{1}{J}\sum_{j=1}^{J}\tilde{R}_{j}\geq R^{\mathrm{th}}.\vspace{-1.5mm}$ (40)
@@
[…]
@@
As a result, the approximated problem can be updated and solved in an iterative manner by existing convex optimization solvers. After solving the subproblem constructed for moving a step ${\mathbf{\Delta b}}_{r}$, we quantify how well the approximate objective function $f_{r}({\mathbf{b}}_{n})=\frac{1}{K}\sum\limits_{k=1}^{K}\left(a_{2}\delta^{2}_{n,k}+a_{3}\delta_{n,k}-g^{\prime}({\mathbf{b}}^{(r)}_{n}){{\mathbf{b}}_{n}}\right)$ predicted the actual reduction in $f$:
$\rho_{r}=\frac{{\mathrm{CRLB}}\bigl{(}{\mathbf{b}}_{n}^{(r)}\bigr{)}\;-\;{\mathrm{CRLB}}\bigl{(}{\mathbf{b}}_{n}^{(r+1)}\bigr{)}}{{\mathrm{CRLB}}\bigl{(}{\mathbf{b}}_{n}^{(r)}\bigr{)}\;-\;f_{r}({\mathbf{b}}_{n}^{(r+1)})},\vspace{-1.5mm}$ (42)
where $\mathrm{CRLB}$ is the true objective (from problem (P3)) evaluated at the current BS position, and $f_{r}({\mathbf{b}}_{n}^{(r+1)})$ is its local approximation in (P3). Then, we have to accept or reject the step. If $\rho_{r}\geq\eta_{s}$, the step is deemed sufficiently accurate and we set ${\mathbf{b}}_{n}^{(r+1)}={\mathbf{b}}_{n}^{(r)}+{\mathbf{\Delta b}}_{r}.$ Otherwise, we reject the step and keep ${\mathbf{b}}_{n}^{(r+1)}={\mathbf{b}}_{n}^{(r)}.$ In addition, we also have to adjust the trust-region radius. Based on $\rho_{r}$, we expand, keep, or shrink the region in which our quadratic model is trusted:
$\epsilon_{r+1}=\begin{cases}\gamma_{n}\,\epsilon_{r},&\rho_{r}\geq\eta_{v},\\ \epsilon_{r},&\eta_{s}\leq\rho_{r}<\eta_{v},\\ \gamma_{d}\,\epsilon_{r},&\rho_{r}<\eta_{s}.\end{cases}\vspace{-1.5mm}$ (43)
@@ V System Performance
Using numerical results in this section, we have studied the fundamental characteristics of ISAC networks and verified the tightness of the expression derived by Monte Carlo simulation results. The system parameters are given based on empirical and modeling studies [37]. Specifically, the number of transmit antennas is $M_{\mathrm{t}}=5$; the number of receive antennas is $M_{\mathrm{r}}=5$; the transmit power is $P_{\mathrm{t}}=0.01$W at each BS; $|\xi|^{2}=10$; the frequency $f^{c}=5$ GHz; the bandwidth $B\in[10,100]$ MHz; the noise power $-120$dB; pathloss coefficients $\alpha=4$ and $\beta=2$. In the following experiments, we consider both one-dimensional and two-dimensional target and user regions. It is important to note that in all scenarios, the BS locations are optimized within a three-dimensional deployment space. In the simulations, we assume that the target and users are located within the same region and are sampled at the same rate.
@@ V-A Convergence and Scaling Law Evaluation
Figure 3: Convergence evaluation of the proposed MM-based algorithm.
@@
(a) BS deployment for optimal communication.
@@
Figure 5: BS deployment comparison for optimal communication and sensing.
@@
Fig. 4 presents the scaling law of the area CRLB as a function of the number of BSs, clearly showing a reduction in CRLB upon increasing the BS number. To ensure a relatively accurate scaling curve, we employ iterative optimization, where at each step we perform a global search to reposition one BS, while holding all others fixed, and repeat this process for each BS until convergence. Moreover, the scaling gain erodes as the sensing-area dimension increases. More explicitly, under the same multiplicative increase in BS number, a 2D region must spread nodes over two orthogonal axes, whereas a 1D region aligns them along a single axis. To validate the periodic-pattern deployment discussed in Remark 2, we took the optimized four-BS layout and replicated it across the entire enlarged region, namely ’simplified constructed solution’ in Fig. 4. The resultant performance closely matches that of the fully optimized scheme, especially in the 1D scenario, demonstrating the practical effectiveness of regular deployments in large-scale networks. Furthermore, in 1D scenarios the theoretical CRLB-scaling bound closely aligns with the simulation results, since the BSs deployed across the $n_{0}$ fundamental unit regions already realize the majority of the localization gain achievable under full BS cooperation. By contrast, in 2D scenarios a persistent gap arises between the theoretical bound and simulation results. The simplified upper bound does not capture the sensing-performance degradation that occurs at the peripheries of individual subregions. By introducing additional BSs that jointly cover the entire 2D area, especially along edges and corners, these boundary effects are substantially alleviated, yielding localization gains beyond those predicted by the bound. Consequently, as the BS density increases, the empirical CRLB scaling in 2D surpasses the theoretical prediction.

### v2

@@ IV-B2 Problem Transformation
$=$ $a_{1}\mathbf{v}^{T}_{n,k}{\mathbf{M}}^{-1}\mathbf{v}_{n,k}\!+\!\text{Tr}\left({\bf{q}}_{1}^{T}\mathbf{v}_{n,k}\right)\!+\!\text{Tr}\left(\mathbf{G}_{n,k}(\mathbf{U}_{n,k}^{c})^{T}{\mathbf{M}}^{-1}\mathbf{U}_{n,k}^{c}\right)\!,\vskip 0.0pt$
where $a_{1}=\mathbf{1}^{T}\mathbf{G}_{n,k}\mathbf{1}+\mathbf{e}_{n}^{T}\mathbf{G}_{n,k}\mathbf{1}+\mathbf{1}^{T}\mathbf{G}_{n,k}\mathbf{e}_{n}+\mathbf{e}^{T}_{n}\mathbf{G}_{n,k}\mathbf{e}_{n}$, and ${\bf{q}}_{1}^{T}=\left({\mathbf{M}}^{-1}\mathbf{U}_{n,k}^{c}\mathbf{G}_{n,k}\mathbf{1}\right)^{T}+\mathbf{1}^{T}\mathbf{G}_{n,k}(\mathbf{U}^{c}_{n,k})^{T}{\mathbf{M}}^{-1}+\mathbf{e}^{T}_{n}\mathbf{G}_{n,k}(\mathbf{U}_{n,k}^{c})^{T}{\mathbf{M}}^{-1}+\left({\mathbf{M}}^{-1}\mathbf{U}_{n,k}^{c}\mathbf{G}_{n,k}{\mathbf{e}}_{n}\right)^{T}$.
@@
$+a_{1}\lambda_{\max}\left({\mathbf{M}}^{-1}\right)+a_{1}\left(\mathbf{v}_{n,k}^{(r)}\right)^{T}{\mathbf{\tilde{M}}}\mathbf{v}_{n,k}^{(r)},\vskip 0.0pt$
@@
where $\tilde{\mathbf{M}}={\mathbf{M}}^{-1}-\lambda_{\max}\left({\mathbf{M}}^{-1}\right){\mathbf{I}}$, and $\lambda_{\max}({\mathbf{M}}^{-1})$ denotes the maximum eigenvalue of matrix ${\mathbf{M}}^{-1}$. In addition, in (34), the term $\mathbf{P}_{n,k}^{T}\mathbf{U}_{n,k}$ can be reduced by discarding any components independent of the optimization variable ${\bf{b}}_{n}$, leaving only the contributions $({{{\mathbf{1}}^{T}}{{\mathbf{P}}_{n,k}^{T}}+{{\mathbf{e}}^{T}_{n}}{{\mathbf{P}}_{n,k}^{T}}})\mathbf{v}_{n,k}$. By omitting the constant terms, the objective function can be approximated as
@@
$\mathop{\mathrm{min}}\limits_{{\mathbf{b}}_{n}}\frac{1}{K}\sum\nolimits_{k=1}^{K}\left(a_{2}\delta^{2}_{n,k}+a_{3}\delta_{n,k}+{\mathbf{q}}_{2}^{T}\frac{\mathbf{t}_{k}-{\mathbf{b}}_{n}}{\left\|\mathbf{t}_{k}-{\mathbf{b}}_{n}\right\|}\right),\vskip 0.0pt$ (37)
@@
where ${\mathbf{q}}_{2}^{T}=2a_{1}\left(\tilde{\mathbf{M}}\mathbf{v}_{n,k}^{(r)}\right)^{T}+{\mathbf{q}}_{1}^{T}+{{{\mathbf{1}}^{T}}{{\mathbf{P}}_{n,k}^{T}}+{{\mathbf{e}}^{T}_{n}}{{\mathbf{P}}_{n,k}^{T}}}$, and $a_{2}={\mathbf{G}}_{n,k}{[n,n]}$, $a_{3}=\frac{1}{2}\sum\nolimits_{m=1,m\neq n}^{N}{\mathbf{G}}_{n,k}{[m,m]}$ $\left\|\mathbf{t}_{k}-{\mathbf{b}}_{m}\right\|^{\beta}$. Here, (37) is still non-convex due to $\frac{\mathbf{t}_{k}-{\mathbf{b}}_{n}}{\left\|\mathbf{t}_{k}-{\mathbf{b}}_{n}\right\|}$. By utilizing the trust region algorithm of [39], we introduce a constraint $\|{\mathbf{b}}_{n}-{\mathbf{b}}_{n}^{(r)}\|\leq\epsilon$ to limit the update region. In the trust region method, a first-order approximation is used at each iteration for constructing a local linear model of the objective function based on its gradient. In practice, this leads to computing the so-called Cauchy point [39], which is obtained by taking a step in the negative gradient’s direction scaled to the boundary of the trust region. Specifically, we adopt Taylor expansion to approximate the objective function as follow.
$\frac{\mathbf{t}_{k}-{\mathbf{b}}_{n}}{\left\|\mathbf{t}_{k}-{\mathbf{b}}_{n}\right\|}\approx\frac{\mathbf{t}_{k}-{\mathbf{b}}_{n}^{(r)}}{\|\mathbf{t}_{k}-{\mathbf{b}}_{n}^{(r)}\|}+g^{\prime}({\mathbf{b}}^{(r)}_{n})({\mathbf{b}}_{n}-{\mathbf{b}}_{n}^{(r)}),\vskip 0.0pt$ (38)
@@
where $g^{\prime}(\mathbf{b}^{(r)}_{n})=-\frac{1}{\|\mathbf{t}_{k}-\mathbf{b}_{n}^{(r)}\|}\,\Biggl(\,\mathbf{I}-\frac{(\mathbf{t}_{k}-\mathbf{b}_{n}^{(r)})(\mathbf{t}_{k}-\mathbf{b}_{n}^{(r)})^{T}}{\|\mathbf{t}_{k}-\mathbf{b}_{n}^{(r)}\|^{2}}\Biggr)$. Following the transformation above, every term in the objective function involving the decision variables is either a convex quadratic or a linear function. Then, for the communication constraint, we introduce auxiliary variables:
$z_{n,j}\geq\|{\mathbf{b}}_{n}-{\mathbf{u}}_{j}\|^{\alpha},\quad\forall j.\vskip 0.0pt$ (39)
@@
We define $C_{j}=1+\frac{1}{\tilde{\sigma}_{c}^{2}}\sum_{m\neq n}g_{m}\|\mathbf{b}_{m}-\mathbf{u}_{j}\|^{-\alpha}$, and apply the first-order Taylor expansion to approximate the communication rate of the user located at ${\bf{u}}_{j}$ as $R_{j}\geq\log_{2}\left(C_{j}+\frac{g_{n}}{\tilde{\sigma}_{c}^{2}z_{n,j}^{(r)}}\right)-\frac{{g_{n}}{}}{\tilde{\sigma}_{c}^{2}(z_{n,j}^{(r)})^{2}\left(C_{j}+\frac{g_{n}}{\tilde{\sigma}_{c}^{2}z_{n,j}^{(r)}}\right)\ln 2}\left(z_{n,j}-z_{n,j}^{(r)}\right)=\tilde{R}_{j}$. Thus, the original non-convex communication constraint can be approximately replaced by the following convex constraint:
$\frac{1}{J}\sum\nolimits_{j=1}^{J}\tilde{R}_{j}\geq R^{\mathrm{th}}.\vskip 0.0pt$ (40)
@@
$(\mathrm{P3}):\quad$ $\min_{{\mathbf{b}}_{n},\delta_{n,k},z_{n,j}}\frac{1}{K}\sum_{k=1}^{K}\left(a_{2}\delta_{n,k}^{2}+a_{3}\delta_{n,k}+\mathbf{q}_{2}^{T}g^{\prime}({\mathbf{b}}_{n}^{(r)}){\mathbf{b}}_{n}\right)$ (41)
@@
As a result, the approximated problem can be updated and solved in an iterative manner by existing convex optimization solvers. After solving the subproblem constructed for moving a step ${\mathbf{\Delta b}}_{r}$, we quantify how well the approximate objective function $f_{r}({\mathbf{b}}_{n})=\frac{1}{K}\sum\nolimits_{k=1}^{K}\left(a_{2}\delta^{2}_{n,k}+a_{3}\delta_{n,k}+\mathbf{q}_{2}^{T}g^{\prime}({\mathbf{b}}^{(r)}_{n}){{\mathbf{b}}_{n}}\right)$ predicted the actual reduction in $f$:
$\rho_{r}=\frac{{\mathrm{CRLB}}\bigl({\mathbf{b}}_{n}^{(r)}\bigr)\;-\;{\mathrm{CRLB}}\bigl({\mathbf{b}}_{n}^{(r+1)}\bigr)}{{\mathrm{CRLB}}\bigl({\mathbf{b}}_{n}^{(r)}\bigr)\;-\;f_{r}({\mathbf{b}}_{n}^{(r+1)})},\vskip 0.0pt$ (42)
where $\mathrm{CRLB}$ is the true objective (from problem (P3)) evaluated at the current BS position, and $f_{r}({\mathbf{b}}_{n}^{(r+1)})$ is its local approximation in (P3). Then, we have to accept or reject the step. If $\rho_{r}\geq\eta_{s}$, the step is deemed sufficiently accurate and we set ${\mathbf{b}}_{n}^{(r+1)}={\mathbf{b}}_{n}^{(r)}+{\mathbf{\Delta b}}_{r}.$ Otherwise, we reject the step and keep ${\mathbf{b}}_{n}^{(r+1)}={\mathbf{b}}_{n}^{(r)}.$ In addition, we also have to adjust the trust‐region radius. Based on $\rho_{r}$, we expand, keep, or shrink the region in which our quadratic model is trusted:
$\epsilon_{r+1}=\begin{cases}\gamma_{n}\,\epsilon_{r},&\rho_{r}\geq\eta_{v},\\ \epsilon_{r},&\eta_{s}\leq\rho_{r}<\eta_{v},\\ \gamma_{d}\,\epsilon_{r},&\rho_{r}<\eta_{s}.\end{cases}\vskip 0.0pt$ (43)
@@
Algorithm 1 Proposed $M$-based Algorithm
1:  Initialize BS positions $\{{\bf{b}}_{n}^{(0)}\}$, sensing target positions $\{{\bf{t}}_{k}\}$, communication user positions $\{{\bf{u}}_{j}\}$, and trust region radius $\epsilon_{0}$.
2:  repeat
3:   for each BS $n=1,\dots,N$ do
4:    Solve problem (P3) to obtain optimal movement $\Delta\mathbf{b}_{n}^{(r)}$ for BS $n$.
5:    Compute predicted and actual reduction ratio $\rho_{r}$ based on (42).
6:    if $\rho_{r}\geq\eta_{s}$ then
7:     Accept update: ${\mathbf{b}}_{n}^{(r+1)}={\mathbf{b}}_{n}^{(r)}+\Delta\mathbf{b}_{n}^{(r)}$
8:    else
9:     Reject update: ${\mathbf{b}}_{n}^{(r+1)}={\mathbf{b}}_{n}^{(r)}$
10:    end if
11:    Update trust-region parameters according to (43)
12:   end for
13:  until convergence of the optimal $\{{\mathbf{b}}_{n}^{*}\}$
@@ V System Performance
Using numerical results in this section, we have studied the fundamental characteristics of ISAC networks and verified the tightness of the expression derived by Monte Carlo simulation results. The system parameters are given based on empirical and modeling studies [40]. Specifically, the number of transmit antennas is $M_{\mathrm{t}}=5$; the number of receive antennas is $M_{\mathrm{r}}=5$; the transmit power is $P_{\mathrm{t}}=0.01$W at each BS; $|\xi|^{2}=10$; the frequency $f^{c}=5$ GHz; the bandwidth $B\in[10,100]$ MHz; the noise power 120-120dB; pathloss coefficients $\alpha=4$ and $\beta=2$. In the following experiments, we consider both one-dimensional and two-dimensional target and user regions. It is important to note that in all scenarios, the BS locations are optimized within a three-dimensional deployment space. In the simulations, we assume that the target and users are located within the same region and are sampled at the same rate.
@@ V-A Convergence and Scaling Law Evaluation
Figure 3: Convergence evaluation of the proposed $M$-based algorithm.
@@
(a) BS deployment for optimal communication achieving $R^{*}$.
@@
(c) BS deployment for maximizing sensing performance with $R_{th}=0.8R^{*}$.
Figure 5: BS deployment under varying sensing-communication weights.
@@
Fig. 4 presents the scaling law of the area CRLB as a function of the number of BSs, clearly showing a reduction in CRLB upon increasing the BS number. To ensure a relatively accurate scaling curve, we employ iterative optimization, where at each step we perform a global search to reposition one BS, while holding all others fixed, and repeat this process for each BS until convergence. Moreover, the scaling gain erodes as the sensing‐area dimension increases. More explicitly, under the same multiplicative increase in BS number, a 2D region must spread nodes over two orthogonal axes, whereas a 1D region aligns them along a single axis. To validate the periodic‐pattern deployment discussed in Remark 2, we took the optimized four-BS layout and replicated it across the entire enlarged region, namely ’simplified constructed solution’ in Fig. 4. The resultant performance closely matches that of the fully optimized scheme, especially in the 1D scenario, demonstrating the practical effectiveness of regular deployments in large-scale networks. Furthermore, in 1D scenarios the theoretical CRLB‐scaling bound closely aligns with the simulation results, since the BSs deployed across the $n_{0}$ fundamental unit regions already realize the majority of the localization gain achievable under full BS cooperation. By contrast, in 2D scenarios a persistent gap arises between the theoretical bound and simulation results. The simplified upper bound does not capture the sensing‐performance degradation that occurs at the peripheries of individual subregions. By introducing additional BSs that jointly cover the entire 2D area, especially along edges and corners, these boundary effects are substantially alleviated, yielding localization gains beyond those predicted by the bound. Consequently, as the BS density increases, the empirical CRLB scaling in 2D surpasses the theoretical prediction.